OpenStudy (lgbasallote):

can someone show me how to use partial fractions in integration please? it confuses me...i think i did well in this during the algebra course...i dont know lol maybe i need a refresher :DDD

OpenStudy (amistre64):

partial fractions is a technique used to rewrite an argument

OpenStudy (amistre64):

its easier to follow if you have a specific example to work thru tho

OpenStudy (unklerhaukus):

\[\frac 1 {x(x-1)}=\frac{A}{x} +\frac B {x-1}\]

OpenStudy (unklerhaukus):

if you were asked to integrate the first expression, it would be easiest to convert to the second expression before integration

OpenStudy (unklerhaukus):

then to solve for A and B

OpenStudy (unklerhaukus):

then to integrate

OpenStudy (lgbasallote):

hmm well how about problems with trig sub AND partial fractions?

OpenStudy (unklerhaukus):

yuk

OpenStudy (lgbasallote):

yuk?

OpenStudy (unklerhaukus):

sorry that sort of problem seams like a nightmare

OpenStudy (lgbasallote):

i know :/

OpenStudy (unklerhaukus):

im sure there are some people who would like that sort of problem but i am not one of them

OpenStudy (lgbasallote):

those guys who gave me brutal problems on the last question might...people like @lalaly @amistre64 and @FoolForMath ? :P

OpenStudy (lalaly):

hi:D

OpenStudy (lgbasallote):

hello :) i'm lgbasallote. nice to meet you lalaly

OpenStudy (lalaly):

lol nice to meet you tooo

OpenStudy (lgbasallote):

enough of the getting-to-know...can you teach me how to solve trig sub + partial fractions combined? you're pretty good in brutal calculus so i thought ill ask you

OpenStudy (anonymous):

lalaly where are you from?

OpenStudy (anonymous):

and how did get the "99"? :p

OpenStudy (lalaly):

lol tai ive been on os for more than a year ... and i solved so many questions :D ... lg id be happy to ...

OpenStudy (anonymous):

partial fractions in integration, is it definite or indefinite?

OpenStudy (anonymous):

remember to + C for the unknown

OpenStudy (anonymous):

and plug in co-ords to find C

OpenStudy (lgbasallote):

yay for lana \m/ indefinite would be nice Tai since definite just comes from indef

OpenStudy (anonymous):

well...just use the power and divide by it?

OpenStudy (anonymous):

let's say x^1/4 becomes (x^5/4)/(5/4) which is 4/5(x^5/4)

OpenStudy (anonymous):

+ C

OpenStudy (lgbasallote):

but that's the power formula..not the partial fractions

OpenStudy (anonymous):

what do you mean by "partial fractions"?

OpenStudy (lgbasallote):

partial fractions,...uhh how are they explained....i dunno lol...

OpenStudy (anonymous):

lol >_>

OpenStudy (anonymous):

the denominator is not factorisable, give me an example to work with plox

OpenStudy (lgbasallote):

that totally depicts my lack of mastery :P anyway...maybe a sample from the legendary @lalaly can hellp us both... \(\Large \int \frac{e^x}{(e^x - 1)(e^x + 3)} dx\)

OpenStudy (anonymous):

lga 1st u have to put e^x=t

OpenStudy (lgbasallote):

uh-huh...

OpenStudy (lgbasallote):

\(\large \int \frac{dt}{(t-1)(t+3)}\)

OpenStudy (anonymous):

eww

OpenStudy (anonymous):

ya gud lga

OpenStudy (lgbasallote):

so it's \(\large \frac{A}{t-1} + \frac{B}{t+3}\)

OpenStudy (lgbasallote):

i dont think this is gonna have trig subs though...

OpenStudy (anonymous):

i never use partail formula ..i use always my trick to solve dis types of question

OpenStudy (lgbasallote):

mind sharing?

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

i'm kinda stuck

OpenStudy (anonymous):

pro people ^_^

OpenStudy (anonymous):

you guys all @ uni?

OpenStudy (lgbasallote):

i think me and lalaly are...gurvinder though i think isnt are you?

OpenStudy (anonymous):

no, i'm lazy

OpenStudy (anonymous):

i'm 17, wbu guys?

OpenStudy (lgbasallote):

hmmm lana's taking a long time..must be some heck of a latex lol

OpenStudy (lgbasallote):

and she's suddenly gone

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

committed suicide

OpenStudy (anonymous):

just kidding

OpenStudy (anonymous):

A/(t−1)+B/(t+3)

OpenStudy (anonymous):

is that the f'(x)?

OpenStudy (anonymous):

or f(x)?

OpenStudy (anonymous):

A(t-1)^-1 + B(t+3)^-1 --> -A(t-1)^-2 -B(t+3)^-2

OpenStudy (unklerhaukus):

@gurvinder what is your trick?

OpenStudy (unklerhaukus):

can you apply to : \[\frac1{x(x−1)}\]

OpenStudy (anonymous):

he's joking, there's no trick

OpenStudy (unklerhaukus):

oh, lame

OpenStudy (anonymous):

yeah...:(

OpenStudy (lgbasallote):

i think it is the cover up method thing

OpenStudy (anonymous):

i wish there were shortcuts to mathematics...

OpenStudy (anonymous):

-drinks coffee-

OpenStudy (lgbasallote):

choose C alway

OpenStudy (turingtest):

@lgbasallote I'd be happy to help you with such an integration problem if you still need the help

OpenStudy (lgbasallote):

well i'll be happy to be helped too @TuringTest after all i still do not know how to it haha lol =))

OpenStudy (anonymous):

ya tia have a trick ok..........i going to tell u all